# SARatrX: Towards Building A Foundation Model for SAR Target Recognition

### v2

@@ A-A Pre-traing dataset setting
The timeline of measured and simulated datasets used for SAR target recognition is illusate in Fig. 6, and we chose the typical open-source dataset as pre-training. As shown in Fig. 7, we collect data from open-source datasets based on our previous research [39, 21]. Now, our pre-training dataset contains 14 open-source SAR target datasets. Here are brief descriptions of each dataset’s targets, scenes, and sensors.
Figure 6: The timeline of measured and simulated datasets used for SAR target recognition (SAR magnitude images were processed in pseudo-color for better visualization). SAR target datasets have increased rapidly since 2020, greatly enriching the diversity of available targets, scenes, and sensors. The different colours represent the target categories included in the dataset. Target categories primarily include fine-grained vehicles, ships, and aircraft. A mixed sample implies the target exhibits more than one category, such as the combination of various vehicles, ships, and aircraft in IRIS-SAR. These scenes may also contain cities, harbors, airports, etc. The sensor devices offer different resolutions and bands for satellite or airborne platforms. However, the number of samples (No. samples) in most datasets is only a few thousand across several categories, due to high costs and the difficulty of annotation. This inspired us to propose SARatrX, with an integrated pre-training dataset SARDet-180K.
Figure 7: Visualization of the fourteen datasets included in our pre-training. In this way, a pre-trained dataset for SAR ATR can be built with multiple targets, scenes, and sensors.
@@
HRSID [63] is a high-resolution dataset for ship detection and instance segmentation based on the European C-band Sentinel-1B, German X-band TerraSAR-X and TanDEM-X satellites. HRSID consists of 5,604 cropped SAR images with 0.5 to 3 m resolutions. The scene is a busy area of maritime transport, such as harbors and estuarine cities, and the annotation targets are ships of different sizes.
@@
MSTAR [18] is the most commonly used target classification dataset released by the Defense Advanced Research Projects Agency, USA. Its sensor is an X-band radar with HH polarization mode and 0.3 m resolution. It contains ten categories of military vehicles with various imaging angles, target variants, and other conditions but with single grass scenes.
@@
SADD [30] is an aircraft detection dataset collected from the German X-band TerraSAR-X satellite. Its resolution is 0.5 m to 3 m with HH polarization. The dataset contains densely parked aircraft of different sizes on airport tarmacs and runways. It has a large number of small-sized planes as well as the airport perimeter area.
@@ A-B Classification dataset for performance Test
We select three target classification datasets, including 25 fine-grained targets from vehicles, ships, aircraft, and others, to evaluate the comprehensive performance of SSL and the foundation model for SAR target recognition. The new SAR classification dataset named SAR-Target in Table VI.
FUSAR-Ship [29] contains 15 primary ship categories and many non-ship targets based on the Gaofen-3 satellite in scenes such as sea, land, coast, river, and island. Based on the experimental setting of [116], we have ten ocean target types, such as four fine-grained ships, bridges, and ocean scene slices.
A-C Hyperparameter settings
TABLE VI: Description of our SAR classification dataset, named SAR-VSA, which contains 25 fine-grained targets. # Train: Number of train samples. # Test: Number of test sample images.
@@ A-C Hyperparameter settings
TABLE VII: Pre-training setting.
@@
TABLE VIII: Classification settings (fine-tuning and linear probing).
@@
TABLE IX: Detailed classification results on MSTAR’s SOCs and EOCs. “-” refers to unknown. Bold indicates the best result, underline is the next best result.
@@
SARatrX 2024 65.3 76.5 76.3 83.4 82.8 85.0 92.1 97.0
TABLE X: Detailed detection results. The metrics are mAP, $\rm{mAP_{50}}$ (@@@@50), $\rm{mAP_{75}}$ (@@@@75), $\rm{mAP_{small}}$ (@@@@s), $\rm{mAP_{medium}}$ (@@@@m), and $\rm{mAP_{large}}$ (@@@@l). “-” refers to unknown. Bold indicates the best result, underline is the next best result.
@@
SARatrX 2024 58.7 86.1 (+5.7) 64.7 (+3.3) 20.0 57.2 49.7
Here are the detailed settings of our pre-training and downstream tasks as shown in Table VII and VIII.
Pre-training - Our default pre-training setting is in Table VII, and other hyperparameters of each method use the default settings from their papers and codes. Our pre-training is applied on 8 NVIDIA RTX3090 GPUs with 200 epochs and 800 batch sizes. Compared to the training settings of MAE, we add ColorJitter ($\rm{contrast}=0.5$) to increase data richness. Moreover, we modify the batch size and epoch according to 8 GPUs. It is worth noting that although MAE uses the normalized pixel value to enhance the feature representation in the visible spectral images, we find that the normalized pixel value cannot be used due to the SAR image noise and prevents the training loss from decreasing properly.
Figure 8: Partial fine-tuning results of HiViT-B and SSL-ImageNet & SAR. Tuning 0 blocks is linear probing. With the increasing number of fine-tuned transformer blocks, our method are consistently better than MAE and experience overfitting later.
Classification setting - All models use the same training settings in downstream classification tasks. Table VIII gives the default setting. Our few-shot learning setting is based on the Dassl toolbox [121, 122] and is averaged over 10 random experiments. Since we focus on the small sample case of the downstream classification task, we use the linear probing method in MAE to finetune models and avoid overfitting.
Partial fine-tuning - Fig. 8 shows why we chose linear probing for few-shot evaluation, as HiViT overfits when many blocks are fine-tuned. Experimental results show that our SAR-KGPA consistently obtains better representations than MAE, and overfitting occurs later.
Appendix B Implementation Details for Section 5 and the Leveraging of SARatrX for Recognition Tasks
@@ B-A Dataset description
We choose MSTAR, the most commonly used dataset in SAR target classification. SOCs are in similar image conditions, and the training set’s depression angle under SOC is 17°, while the test set is 15°. Ten categories of targets include BMP2, BRDM2, BTR60, BTR70, T62, T72, 2S1, D7, ZIL131, and ZSU234 as shown in Table VI. The EOCs setting is the imaging condition variations to test robustness. Existing methods have been saturated with different experimental settings of MSTAR [25], but few samples and depression angle variations remain challenging. We use SARDet-100K and OGSOD datasets, which have many test samples and categories, to evaluate the detection performance fully. The OGSOD comparison results are derived from the original article’s [22] single-modal approach using only SAR images. SSDD and SAR-Aircraft are ship and aircraft categories.
@@ B-B Hyperparameter settings
Classification setting is follow Table VIII. The only difference is that we use partial fine-tuning for better performance, and the last 6 blocks are added to the fine-tuned.
Detection setting is follows the default setting in HiViT, and we adjust the learning rate to 5e-4. We used the same settings for each dataset fine-tuning, see our GitHub configuration based on the mmdetection [123] framework for details.
@@ B-C Detailed results
We provide detail classification and detection result on Table IX and X. Although the proposed method outperforms existing methods in mAP, there is still scope for improvement in some refined metrics.

### v3

@@ A-A Pre-traing dataset setting
we chose the most open-source dataset as pre-training. As shown in Fig. 8, we collect data from open-source datasets based on our previous research [37, 39]. Now, our pre-training dataset contains 14 open-source SAR target datasets. Here are brief descriptions of each dataset’s targets, scenes, and sensors.
Figure 9: Timeline of measured and simulated datasets used for SAR target recognition (SAR magnitude images were processed in pseudo-color for better visualization). SAR target datasets have increased rapidly since 2020, greatly enriching the diversity of available targets, scenes, and sensors. The different colours represent the target categories included in the dataset. Target categories primarily include fine-grained vehicles, ships, and aircraft. A mixed sample implies the target exhibits more than one category, such as the combination of various vehicles, ships, and aircraft in IRIS-SAR. These scenes may also contain cities, harbors, airports, etc. The sensor devices offer different resolutions and bands for satellite or airborne platforms. However, the number of samples (No. samples) in most datasets is only a few thousand across several categories, due to high costs and the difficulty of annotation. This inspired us to propose SARATR-X, with an integrated pre-training dataset SARDet-180K.
MSTAR [36] is the most commonly used target classification dataset released by the Defense Advanced Research Projects Agency, USA. Its sensor is an X-band radar with HH polarization mode and 0.3 m resolution. It contains ten categories of military vehicles with various imaging angles, target variants, and other conditions but with simple grass scenes and not completely open source.
@@
HRSID [82] is a high-resolution dataset for ship detection and instance segmentation based on the European C-band Sentinel-1B, German X-band TerraSAR-X and TanDEM-X satellites. HRSID consists of 5,604 cropped SAR images with 0.5 to 3 m resolutions. The scene is a busy area of maritime transport, such as harbors and estuarine cities, and the annotation targets are civil ships of different sizes.
SADD [23] is an aircraft detection dataset collected from the German X-band TerraSAR-X satellite. Its resolution is 0.5 m to 3 m with HH polarization. The dataset contains densely parked aircraft of different sizes on airport tarmacs, runways, and airport cemeteries. It has a large number of small-sized planes as well as the airport perimeter area.
@@ A-B Classification dataset for performance Test
TABLE VIII: Description of our SAR classification dataset, named SAR-VSA, which contains 25 target classes. # Train: Number of train samples. # Test: Number of test samples.
@@
TABLE IX: Pre-training setting.
@@
TABLE X: Classification settings (fine-tuning and linear probing).
@@
Figure 10: Partial fine-tuning results of HiViT-B and SSL-ImageNet & SAR. Tuning 0 blocks is linear probing. With the increasing number of fine-tuned transformer blocks, our method is consistently better than MAE and experiences overfitting later under the 5-shot setting.
TABLE XI: Detailed classification results on MSTAR’s SOCs and EOCs. “-” refers to unknown. Bold indicates the best result, underline is the next best result.
@@
SARATR-X 2024 65.3 76.5 76.3 83.4 82.8 85.0 92.1 97.0
We select three target classification datasets, including 25 fine-grained targets from vehicles, ships, aircraft, and others, to evaluate the comprehensive performance of SSL and the foundation model for SAR target recognition. The new SAR classification dataset is named SAR-Target in Table VIII.
FUSAR-Ship [69] contains 15 primary ship categories and many non-ship targets based on the Gaofen-3 satellite in scenes such as sea, land, coast, river, and island. Based on the experimental setting of [128], we have ten ocean target types, such as four fine-grained ships, bridges, ocean scenes, and false alarm slices.
A-C Hyperparameter settings
TABLE XII: Detailed detection results. The metrics are mAP, $\rm{mAP_{50}}$ (@@@@50), $\rm{mAP_{75}}$ (@@@@75), $\rm{mAP_{small}}$ (@@@@s), $\rm{mAP_{medium}}$ (@@@@m), and $\rm{mAP_{large}}$ (@@@@l). “-” refers to unknown. Bold indicates the best result, underline is the next best result.
@@ A-C Hyperparameter settings
SARATR-X 2024 58.7 86.1 (+5.7) 64.7 (+3.3) 20.0 57.2 49.7
Here are the detailed settings of our pre-training and downstream tasks as shown in Table IX and X.
Pre-training - Our default pre-training setting is in Table IX, and other hyperparameters of each method use the default settings from their papers and codes. Our pre-training is applied on 8 NVIDIA RTX3090 GPUs with 200 epochs and 800 batch sizes. Compared to the training settings of MAE, we add ColorJitter ($\rm{contrast}=0.5$) to increase data richness. Moreover, we modify the batch size and epoch according to 8 GPUs. It is worth noting that although MAE uses the normalized pixel value to enhance the feature representation in the visible spectral images, we find that the normalized pixel value cannot be used due to the SAR image noise and prevents the training loss from decreasing properly.
Classification setting - All models use the same training settings in downstream classification tasks. Table X gives the default setting. Our few-shot learning setting is based on the Dassl toolbox [129, 130] and is averaged over 10 random experiments. Since we focus on the small sample case of the downstream classification task, we use the linear probing method in MAE to finetune models and avoid overfitting.
Partial fine-tuning - Fig. 10 shows why we chose linear probing for few-shot evaluation, as HiViT overfits when many blocks are fine-tuned. Experimental results show that our method consistently obtains better representations than MAE, and overfitting occurs later.
Appendix B Implementation Details for Section 5 and the Leveraging of SARATR-X for Recognition Tasks
@@ B-A Dataset description
We choose MSTAR, the most commonly used dataset in SAR target classification. SOCs are in similar image conditions, and the training set’s depression angle under SOC is 17°, while the test set is 15°. Ten categories of targets include BMP2, BRDM2, BTR60, BTR70, T62, T72, 2S1, D7, ZIL131, and ZSU234 as shown in Table VIII. EOCs settings are the imaging condition variations to test robustness. Existing methods have been saturated with different experimental settings of MSTAR [18], but few samples and depression angle variations remain challenging.
We use SARDet-100K and OGSOD datasets, which have many test samples and categories, to evaluate the detection performance fully. The OGSOD comparison results are derived from the original article’s [45] single-modal approach using only SAR images. SSDD and SAR-Aircraft are ship and aircraft categories.
@@ B-B Hyperparameter settings
Classification setting is follow Table X. The only difference is that we use partial fine-tuning for better performance, and the last 6 blocks are used to the fine-tuned.
Detection setting follows the default setting in HiViT, and we adjust the learning rate to 5e-4. We used the same settings for each dataset fine-tuning, and our GitHub configuration is based on the mmdetection [131] framework for details.
@@ B-C Detailed results
We provide detailed classification and detection results in Table XI and XII. Although the proposed method outperforms existing methods in mAP, there is still scope for improvement in some refined metrics.